# CogGuide: Human-Like Guidance for Zero-Shot Omni-Modal Reasoning

### v2

@@ 3.2 Main Results
Tables 3–5 report detailed results on the IntentBench, WorldSense, and Daily-Omni datasets, respectively. The “three-module” scheme consistently outperforms its respective baselines across all combinations of the four Pipelines and the three reasoning engines. The best results are sometimes achieved by CG_Qwen_vl (using Qwen2.5-VL-32B as the intent perceiver) and sometimes by the CG_GLM_vl (using GLM-4.5V as the intent perceiver), but both surpass “no intent” (Abl_NoIntent) and “single strategy” (Abl_SingleStrategy). In addition, both Abl_NoIntent and Abl_SingleStrategy outperform the baseline, confirming the effectiveness of each module.
@@
Pipeline Reasoning Model (Baseline) CG_Qwen _vl CG_GLM _vl Abl_ NoIntent Abl_ Single Strategy
@@
Pipeline Reasoning Model (Baseline) CG_Qwen _vl CG_GLM _vl Abl_ NoIntent Abl_ Single Strategy
@@
Pipeline Reasoning Model (Baseline) CG_Qwen _vl CG_GLM _vl Abl_ NoIntent Abl_ Single Strategy
@@
From Table 3, on the IntentBench dataset: for HumanOmniV2, the baseline accuracy is 69.33%, and with our method it reaches a maximum of 71.18% (+1.85 pp; CG_Qwen_vl: Qwen2.5-VL-32B intent perceiver + 3 strategies + selector; Pipeline = Qwen3). For Qwen2.5-Omni, the baseline is 64.20%, with a maximum of 66.07% (+1.87 pp; CG_GLM_vl: GLM-4.5V intent perceiver + 3 strategies + selector; Pipeline = GPT-4o). For Qwen2.5-VL, the baseline is 61.68%, with a maximum of 63.83% (+2.15 pp; CG_GLM_vl; Pipeline = Qwen3).
On the WorldSense dataset (Table 4): HumanOmniV2 baseline 47.10%, maximum 48.80% (+1.70 pp; CG_Qwen_vl; Pipeline = GPT-4o); Qwen2.5-Omni baseline 45.40%, maximum 47.86% (+2.46 pp; CG_Qwen_vl; Pipeline = Qwen3); Qwen2.5-VL baseline 37.39%, maximum 43.41% (+6.02 pp; CG_Qwen_vl; Pipeline = GLM-4.5).
Similarly, on the Daily-Omni dataset (Table 5): HumanOmniV2 baseline 58.47%, maximum 62.74% (+4.27 pp; CG_Qwen_vl; Pipeline = GLM-4.5); Qwen2.5-Omni baseline 47.45%, maximum 56.96% (+9.51 pp; CG_GLM_vl; Pipeline = Qwen3); Qwen2.5-VL baseline 47.28%, maximum 51.71% (+4.43 pp; CG_GLM_vl; Pipeline = GLM-4.5).
In summary, the complete “three-module” intent-sketch reasoning scheme achieves stable improvements across all experiments (with a maximum gain of approximately 9.51%). Notably, these gains do not depend on a specific Pipeline model: even when the strategy generation and selection modules are replaced with smaller open-source models, large-model reasoning performance still exceeds the baseline without strategy prompts. This indicates that our method offers strong portability and plug-and-play characteristics, delivering consistent performance gains by embedding a unified strategy-prompting process regardless of the underlying model or platform interface.
@@ 3.3 Ablation Study and Analyses
To further understand the mechanisms of the three modules, we conduct ablations on Intent Perceiver, Strategy Generator, Strategy Selector. As shown in Tables 3–5, removing any single module leads to a drop in accuracy: taking Daily-Omni as an example, the complete HumanOmniV2 × GLM-4.5 configuration achieves 62.74%; removing the Intent Perceiver module (Abl_NoIntent) reduces it to 59.31% (−3.43 pp), while switching to a single strategy (Abl_SingleStrategy) yields 60.74% (−2.00 pp). Similar trends hold for other combinations: on WorldSense with Qwen2.5-VL × GLM-4.5, the complete configuration reaches 43.41%, while “no Intent Perceiver” gives 42.40% (−1.01 pp) and “single strategy” yields 41.93% (−1.48 pp); on IntentBench with HumanOmniV2 × Qwen3, the complete setting attains 71.18%, “no Intent Perceiver” 70.82% (−0.36 pp), and “single strategy” 69.96% (−1.22 pp). Without the intent provided by the Intent Perceiver, the model may miss crucial information, typically resulting in a modest performance drop; and when bypassing Strategy Selector to use a single strategy directly, the reasoning path may be suboptimal, leading to a more pronounced decline in accuracy. Notably, under some strong Pipeline configurations, the contribution of the Intent Perceiver module is relatively small (e.g., IntentBench with Qwen2.5-Omni × GPT-4o: complete 65.95% vs. no intent 65.82%, +0.13 pp), possibly because such powerful generative models already possess certain scene-understanding capabilities, partially offsetting the benefits of additional intent. Overall, “multi-strategy + selector” contributes the primary gains, while Intent Perceiver provides more substantial additional improvements in scenarios with stronger video/audio dependence (e.g., Daily-Omni). These ablation results indicate that the three-module component we propose performs distinct roles and that their synergy is necessary to maximally enhance reasoning performance: the Intent Perceiver supplies a clear scene context, the Strategy Generator broadens the solution-space of reasoning trajectories, and the Strategy Selector ensures the choice of the optimal approach, thereby making the final answers more accurate and reliable.
Our method does not depend on a specific large model or inference framework. As seen from the vertical comparisons in Tables 3–5, it consistently yields positive improvements across different reasoning engines (HumanOmniV2, Qwen2.5-Omni, Qwen2.5-VL) and different Pipelines (GPT-4o / GLM-4.5 / Doubao-Seed-1.6 / Qwen3). Moreover, for base models without any post-training, the relative gains from strategy prompting are more pronounced: for example, on WorldSense, Qwen2.5-VL improves from 37.39% to 43.41% (+6.02 pp; CG_Qwen_vl; Pipeline = GLM-4.5), and on HumanOmniV2 from 47.10% to 48.80% (+1.70 pp; CG_Qwen_vl; Pipeline = GPT-4o). This indicates that base models lacking post-training do not necessarily possess robust intent extraction and sound reasoning capabilities, and therefore benefit more from additional intent and strategy information. From the perspective of the Pipeline modules, the models we use span different “platforms,” including closed-source APIs and open-source models. The results show that regardless of which Pipeline model is adopted, the three-module strategy prompting has a positive effect on each reasoning engine: for example, on Daily-Omni, HumanOmniV2 improves from 58.47% to 62.74% (+4.27 pp; CG_Qwen_vl; Pipeline = GLM-4.5); even when replacing the Pipeline from GLM-4.5 (355B/32B) with the much smaller Qwen3 (235B/22B), the improvement for HumanOmniV2 remains close to four percentage points (+4.02 pp). This demonstrates that our method has strong platform-agnostic adaptability and generality: as a reasoning enhancement independent of underlying models, strategy prompting can be seamlessly embedded into different large-model inference engines to deliver cross-model performance gains.
Taken together, these findings show that the intent recognition, multi-strategy generation, and selection modules each contribute gains; the improvements do not rely on any particular reasoning engine or Pipeline; they are compatible with both closed- and open-source models of varying sizes; and the component operates in a plug-and-play manner, maintaining stable improvements across open-/closed-source settings and model scales, and under different engine/front-end substitutions—demonstrating cross-model, cross-platform, and cross-scenario generality and robustness.
@@ CONCLUSION
This paper proposes a “human-like cognitive intent sketch” component that explicitly models the “understand–plan–select” reasoning process via a three-module pipeline (Intent Perceiver, Strategy Generator, Strategy Selector). As a plug-and-play, context-injected strategy, it improves the controllability and reliability of multimodal reasoning under a zero-shot setting without parameter fine-tuning, and offers strong portability and practicality across modalities, reasoning engines, and platforms. From an information-theoretic perspective, “multi-strategy generation + strategy selection” produces complementary candidates conditioned on the intent representation and selects the best by a minimum conditional-entropy/Bayesian-risk criterion, thereby effectively contracting the posterior over answers and reducing decision uncertainty and empirical error; this mechanism explains the stable gains and stronger interpretability achieved without any training overhead.
Experimental results show that on the IntentBench, WorldSense, and Daily-Omni benchmarks, the complete “three-module” scheme outperforms its respective baselines across all combinations of Pipelines and reasoning engines, with a maximum improvement of about +9.51 percentage points. Ablation analyses further indicate that the three modules are complementary and synergistic: “multi-strategy + selector” constitutes the primary source of gains, while explicit intent contributes especially in tasks with stronger video/audio dependence. Meanwhile, the method yields stable, cross-platform performance improvements for various reasoning engines and diverse Pipeline models, reflecting its plug-and-play operation and strong transferability. In sum, the “intent sketch” reasoning component provides a lightweight yet effective paradigm for enhancing the alignment, robustness, and interpretability of complex cross-modal reasoning.

### v3

@@ 3.2 Main Results
Pipeline Reasoning Model (Baseline) CG_Qwen CG_GLM Abl_NI Abl_SP
@@
Pipeline Reasoning Model (Baseline) CG_Qwen CG_GLM Abl_NI Abl_SP
@@
Pipeline Reasoning Model (Baseline) CG_Qwen CG_GLM Abl_NI Abl_SP
@@
From Table 3, on the IntentBench dataset: for HumanOmniV2, the baseline accuracy is 69.33%, and with our method it reaches a maximum of 71.18% (+1.85 pp; CG_Qwen; Pipeline = Qwen3). For Qwen2.5-Omni, the baseline is 64.20%, with a maximum of 66.07% (+1.87 pp; CG_GLM; Pipeline = GPT-4o). For Qwen2.5-VL, the baseline is 61.68%, with a maximum of 63.83% (+2.15 pp; CG_GLM; Pipeline = Qwen3).
On the WorldSense dataset (Table 4): HumanOmniV2 baseline 47.10%, maximum 48.80% (+1.70 pp; CG_Qwen; Pipeline = GPT-4o); Qwen2.5-Omni baseline 45.40%, maximum 47.86% (+2.46 pp; CG_Qwen; Pipeline = Qwen3); Qwen2.5-VL baseline 37.39%, maximum 43.41% (+6.02 pp; CG_Qwen; Pipeline = GLM-4.5).
Similarly, on the Daily-Omni dataset (Table 5): HumanOmniV2 baseline 58.47%, maximum 62.74% (+4.27 pp; CG_Qwen; Pipeline = GLM-4.5); Qwen2.5-Omni baseline 47.45%, maximum 56.96% (+9.51 pp; CG_GLM; Pipeline = Qwen3); Qwen2.5-VL baseline 47.28%, maximum 51.71% (+4.43 pp; CG_GLM; Pipeline = GLM-4.5).
Table 3-5 report detailed results on three benchmarks, respectively. The best results are sometimes achieved by CG_Qwen and sometimes by the CG_GLM, but both surpass “no intent” (Abl_NI) and “single policy” (Abl_SP). In addition, both Abl_NI and Abl_SP outperform the baseline, confirming the effectiveness of each module. The maximum gain corresponds to +9.51 pp (a 20.04% relative improvement). Importantly, the three-module scheme consistently outperforms the corresponding baselines across all combinations of the four pipelines and three reasoning engines, regardless of whether the pipeline LLMs are closed-source or open-source, this indicates that our method offers strong portability and plug-and-play characteristics.
@@ 3.3 Ablation Study and Analyses
We ablate the three modules—Intent Perceiver (IP), Policy Generator (PG), and Strategy Selector (SS)—with results summarized in Table III. Removing any single component degrades accuracy. For clarity, we denote Abl_NI as removing IP (i.e., no $Z_{\mathrm{IP}}$), and Abl_SP as using a single policy (N=1N{=}1), which disables PG’s diversity and makes SS degenerate to selecting the only sketch. For Daily-Omni (HumanOmniV2 $\times$ GLM-4.5), the full system reaches $62.74\%$; Abl_NI drops to $59.31\%$ (3.43-3.43 pp), and Abl_SP to $60.74\%$ (2.00-2.00 pp). Similar patterns hold elsewhere: on WorldSense (Qwen2.5-VL $\times$ GLM-4.5) full $43.41\%$, no-IP $42.40\%$ (1.01-1.01 pp), single-policy $41.93\%$ (1.48-1.48 pp); on IntentBench (HumanOmniV2 $\times$ Qwen3) full $71.18\%$, no-IP $70.82\%$ (0.36-0.36 pp), single-policy $69.96\%$ (1.22-1.22 $p$).
Interpretation. Without the intent provided by the Intent Perceiver, the model may miss crucial omni-modal cues, typically resulting in a modest performance drop; with Abl_SP (N=1N{=}1), the strategy set collapses and coverage over semantic classes $\mathcal{M}$ shrinks, which aligns with the increase of semantic uncertainty in Eqs. (2)–(3). Note that Abl_SP does not remove SS; rather, the selector trivially chooses the only available sketch. Therefore, the observed decline under Abl_SP should be attributed to the loss of PG-induced diversity (and the $\mathrm{Div}(\cdot)$ term in Eq. (5)). Among the ablations, the impact of policy generation is larger: the single-policy (Abl_SP) degradation tends to exceed the no-IP case, indicating that multi-path thinking is intrinsically valuable. This mirrors the human “understand–plan–select” pattern, where careful multi-path simulation precedes a final decision. The selector’s role is captured by minimum conditional entropy/Bayesian risk (Eqs. (6)–(9)); when N>1N{>}1 and SS actively scores candidate sketches, uncertainty is further reduced in line with the contraction in Eqs. (11)–(13). From the pipeline perspective, Qwen3-based pipelines exhibit the highest win rate across settings; given that most reasoning engines here are from the Qwen family, this “same-lineage” pairing likely improves adherence to strategy guidance and reduces mismatch between policy sketches and the executor. From the reasoning-model perspective, omni models outperform VL models, suggesting that access to audio cues materially improves disambiguation and intent grounding.
Our approach is model- and framework-agnostic. Across reasoning engines (HumanOmniV2, Qwen2.5-Omni, Qwen2.5-VL) and pipelines (GPT-4o / GLM-4.5 / Doubao-Seed-1.6 / Qwen3), improvements are consistent. Gains are larger for base models lacking post-training (e.g., WorldSense: Qwen2.5-VL $37.39\%\rightarrow 43.41\%$, +6.02+6.02 pp, CG_Qwen, Pipeline=GLM-4.5; HumanOmniV2 $47.10\%\rightarrow 48.80\%$, +1.70+1.70 pp, Pipeline=GPT-4o). On Daily-Omni, HumanOmniV2 improves $58.47\%\rightarrow 62.74\%$ (+4.27+4.27 pp) with GLM-4.5; replacing GLM-4.5 with the smaller Qwen3 still yields +4.02+4.02 pp. These results indicate a plug-and-play component that reliably enhances reasoning across open/closed models and scales, demonstrating cross-model, cross-platform, and cross-scenario robustness.
@@ CONCLUSION
This paper presents a zero-shot omni-modal reasoning component that implements a plug-and-play, intent-sketch pipeline: Intent Perceiver, Policy Generator, and Strategy Selector, which uses context injection to enhance reasoning performance without fine-tuning and generalizes across modalities, reasoning engines, and platforms. From an information-theoretic view, Policy Generator produces complementary candidates conditioned on intent, Strategy Selector selects among candidate strategies via prompt-based evaluation under additional task-specific conditioning, effectively contracting the answer posterior and reducing decision uncertainty, explaining the stable gains and improved interpretability without training overhead.
On IntentBench, WorldSense, and Daily-Omni, the full three-module scheme consistently outperforms baselines across all Pipeline/engine combinations, with maximum gains of +9.51 pp (relative improvement of 20.04%). Ablations show complementary roles: PG+SS is the primary source of improvement, while explicit intent is especially beneficial for video/audio-dependent tasks. The method delivers stable cross-platform gains for diverse reasoning engines and Pipeline models, underscoring its plug-and-play nature and strong transferability. Overall, the intent sketch offers a lightweight, effective paradigm for improving the alignment, robustness, and interpretability of complex cross-modal reasoning.